### v3

@@ 2.4 Measurability Issues
To address these measurability issues, we use the concept of outer measure $\mathbb{P}^{*}$ and inner measure $\mathbb{P}_{*}$ frequently adopted in empirical process theory [31, Section 1.2-5]. Given a probability space $(\Xi,\mathscr{G},\mathbb{P})$, every subset $B\subseteq\Xi$ is assigned with an outer measure:
$\mathbb{P}^{*}(B)=\inf\{\mathbb{P}(A):B\subseteq A,A\in\mathscr{G}\}.$
Notably, $\mathbb{P}^{*}$ is subadditive: $\mathbb{P}^{*}(B_{1}\cup B_{2})\leq\mathbb{P}^{*}(B_{1})+\mathbb{P}^{*}(B_{2})$, following the union bound. Correspondingly, the inner measure is $\mathbb{P}_{*}(B)=\sup\{\mathbb{P}(A):B\supseteq A,A\in\mathscr{G}\}$. Clearly, $\mathbb{P}_{*}(B)=1-\mathbb{P}^{*}(\Xi\backslash B)$. Finally, when $B\in\mathscr{G}$ is measurable, then the inner and outer measure agree: $\mathbb{P}^{*}(B)=\mathbb{P}_{*}(B)=\mathbb{P}(B)$.
@@ 3.1 Uniform Subdifferential Bounds via Selections
Our proof is structured into two parts. In the first part, we demonstrate that Theorem 1 holds under the assumption that $\mathcal{O}$ is open and convex, and that both $f_{1}$ and $f_{2}$ are convex on $\mathcal{O}$.
In the second part, we address the general case where $f_{1}$ and $f_{2}$ are locally weakly convex on an open set $\mathcal{O}$. We show that this scenario can be reduced to the convex case, thus concluding our proof of Theorem 1.
Part I. In this part, we address the case where $f_{1}$ and $f_{2}$ are both real-valued convex functions on an open and convex set $\mathcal{O}$. Recall the definition of Hausdorff distance:
$\mathbb{H}(\partial f_{1}(x),\partial f_{2}(x))=\max\{\mathbb{D}(\partial f_{1}(x),\partial f_{2}(x)),\mathbb{D}(\partial f_{2}(x),\partial f_{1}(x))\}.$
By symmetry, it suffices to prove $\sup_{x\in\mathcal{O}}\mathbb{D}(\partial f_{1}(x),\partial f_{2}(x))\leq\sup_{x\in\mathcal{O}}\left\|{g_{1}(x)-g_{2}(x)}\right\|$, or equivalently,
$\sup_{x\in\mathcal{O}}\sup_{y_{1}\in\partial f_{1}(x)}\inf_{y_{2}\in\partial f_{2}(x)}\left\|{y_{1}-y_{2}}\right\|\leq\sup_{x\in\mathcal{O}}\left\|{g_{1}(x)-g_{2}(x)}\right\|.$ (3.2)
By applying convex duality, we obtain:
$\begin{split}\inf_{y_{2}\in\partial f_{2}(x)}\left\|{y_{1}-y_{2}}\right\|&=\inf_{y_{2}\in\partial f_{2}(x)}\sup_{v:\left\|{v}\right\|\leq 1}\langle y_{1}-y_{2},v\rangle=\sup_{v:\left\|{v}\right\|\leq 1}\inf_{y_{2}\in\partial f_{2}(x)}\langle y_{1}-y_{2},v\rangle,\end{split}$ (3.3)
where the interchangeability between inf and sup in the second identity is due to Sion’s minimax theorem [33]. By substituting this identity back into equation (3.2), our goal reduces to prove:
$\sup_{x\in\mathcal{O}}\sup_{y_{1}\in\partial f_{1}(x)}\sup_{v:\left\|{v}\right\|\leq 1}\inf_{y_{2}\in\partial f_{2}(x)}\langle y_{1}-y_{2},v\rangle\leq\sup_{x\in\mathcal{O}}\left\|{g_{1}(x)-g_{2}(x)}\right\|.$
This ultimately reduces to proving for any $x\in\mathcal{O}$ and $y_{1}\in\partial f_{1}(x)$, and any $v$ with $\left\|{v}\right\|\leq 1$:
$\inf_{y_{2}\in\partial f_{2}(x)}\langle y_{1}-y_{2},v\rangle\leq\sup_{x\in\mathcal{O}}\left\|{g_{1}(x)-g_{2}(x)}\right\|.$ (3.4)
To prove this inequality, we invoke a perturbation argument standard in convex analysis. We start by taking a sequence $t_{n}\to 0$ as $n\to\infty$ with $t_{n}>0$ and $x+t_{n}v\in\mathcal{O}$ per $n\in\mathbb{N}$. Such a sequence $t_{n}$ exists because $\mathcal{O}$ is open. Next, we pick any cluster point $\tilde{y}_{2}$ of the sequence $\{g_{2}(x+t_{n}v)\}_{n\in\mathbb{N}}$. Given that $g_{2}(x+t_{n}v)\in\partial f_{2}(x+t_{n}v)$, and the subdifferential map $z\mapsto\partial f_{2}(z)$ is locally bounded [34, Proposition 6.2.2], a cluster point $\tilde{y}_{2}$ necessarily exists in $\mathbb{R}^{d}$. Furthermore, because the subdifferential map $z\mapsto\partial f_{2}(z)$ is outer semicontinuous [34, Theorem 6.2.4], $\tilde{y}_{2}$ must belong to the set $\partial f_{2}(x)$. Finally, by selecting a subsequence of $t_{n}$ if necessary, we can assume W.L.O.G. that the following limit holds:
$\lim_{n\to\infty}g_{2}(x+t_{n}v)=\tilde{y}_{2}.$
Given convexity of $f_{1}$, $y_{1}\in\partial f_{1}(x)$ and $g_{1}(x+t_{n}v)\in\partial f_{1}(x+t_{n}v)$, we obtain for every $n\in\mathbb{N}$:
$\langle y_{1},v\rangle\leq\frac{1}{t_{n}}(f_{1}(x+t_{n}v)-f_{1}(x))\leq\langle g_{1}(x+t_{n}v),v\rangle.$
Thus, we have for every $v$ with $\left\|{v}\right\|\leq 1$:
$\begin{split}\inf_{y_{2}\in\partial f_{2}(x)}\langle y_{1}-y_{2},v\rangle&\leq\langle y_{1}-\tilde{y}_{2},v\rangle\leq\liminf_{n\to\infty}\langle g_{1}(x+t_{n}v)-g_{2}(x+t_{n}v),v\rangle,\end{split}$
which by an application of Cauchy-Schwartz inequality to the RHS, yields:
$\inf_{y_{2}\in\partial f_{2}(x)}\langle y_{1}-y_{2},v\rangle\leq\liminf_{n\to\infty}\left\|{g_{1}(x+t_{n}v)-g_{2}(x+t_{n}v)}\right\|\leq\sup_{x\in\mathcal{O}}\left\|{g_{1}(x)-g_{2}(x)}\right\|.$
This shows that the desired inequality (3.4) holds for any $x\in\mathcal{O}$ and $y_{1}\in\partial f_{1}(x)$, and any $v$ with $\left\|{v}\right\|\leq 1$. Thus, we successfully establish the assertion of Theorem 1 in the convex settings.
Part II. In the second part, we study the general case where $f_{1}$ and $f_{2}$ are both locally weakly convex on an open set $\mathcal{O}$, extending the established results for convex functions in Part I.
Let $z\in\mathcal{O}$ be arbitrary. Then there is a convex, open neighborhood of $U_{z}\subseteq\mathcal{O}$ such that $f_{1}$ and $f_{2}$ are weakly convex on $U_{z}$. To utilize the established results for convex functions in Part I, we transform $f_{1}$ and $f_{2}$ into convex functions on $U_{z}$ by adding a quadratic regularization. Specifically, we consider:
$\tilde{f}_{1}=f_{1}+\frac{L}{2}\left\|{\cdot}\right\|^{2}~{}~{}\text{and}~{}~{}\tilde{f}_{2}=f_{2}+\frac{L}{2}\left\|{\cdot}\right\|^{2}.$
where $L$ is a large constant. This addition renders both $\tilde{f}_{1}$ and $\tilde{f}_{2}$ convex on $U_{z}$.
As a result of this transformation, we modify the selections from the subdifferentials accordingly:
$\tilde{g}_{1}(x)=g_{1}(x)+Lx~{}~{}\text{and}~{}~{}\tilde{g}_{2}(x)=g_{2}(x)+Lx.$
Importantly, this modification does not affect the Hausdorff distance between the subdifferentials nor the norm of the difference between selections from the subdifferentials, i.e., for every $x\in U_{z}$:
$\mathbb{H}(\partial f_{1}(x),\partial f_{2}(x))=\mathbb{H}(\partial\tilde{f}_{1}(x),\partial\tilde{f}_{2}(x))~{}~{}\text{and}~{}~{}\left\|{g_{1}(x)-g_{2}(x)}\right\|=\left\|{\tilde{g}_{1}(x)-\tilde{g}_{2}(x)}\right\|.$
Since $\tilde{f}_{1}$ and $\tilde{f}_{2}$ are convex on the open, convex set $U_{z}$, and both $\tilde{g}_{1}$ and $\tilde{g}_{2}$ are selections from $\partial\tilde{f}_{1}$ and $\partial\tilde{f}_{2}$, we can apply the result as proved in Part I to these transformed functions to conclude:
$\begin{split}\sup_{x\in U_{z}}\mathbb{H}(\partial f_{1}(x),\partial f_{2}(x))&=\sup_{x\in U_{z}}\mathbb{H}(\partial\tilde{f}_{1}(x),\partial\tilde{f}_{2}(x))\\ &\leq\sup_{x\in U_{z}}\left\|{\tilde{g}_{1}(x)-\tilde{g}_{2}(x)}\right\|=\sup_{x\in U_{z}}\left\|{g_{1}(x)-g_{2}(x)}\right\|.\end{split}$
Since $z\in U_{z}$, and $U_{z}\subseteq\mathcal{O}$, it follows that for every $z\in\mathcal{O}$:
$\mathbb{H}(\partial f_{1}(z),\partial f_{2}(z))\leq\sup_{x\in\mathcal{O}}\left\|{g_{1}(x)-g_{2}(x)}\right\|.$
Taking supremum over $z\in\mathcal{O}$, this observation directly leads to the conclusion of Theorem 1:
$\sup_{x\in\mathcal{O}}\mathbb{H}(\partial f_{1}(x),\partial f_{2}(x))\leq\sup_{x\in\mathcal{O}}\left\|{g_{1}(x)-g_{2}(x)}\right\|.$
@@
Let $\mathcal{O}\subseteq\mathbb{R}^{d}$ be open. For a pair of locally weakly convex functions $f_{i}:\mathcal{O}\to\mathbb{R}$ for $i=1,2$, we define:
@@
In defining ${\rm d}_{2}$, for a locally weakly convex $f:\mathcal{O}\to\mathbb{R}$, we recall the graph of its subdifferential:
@@
where the Cartesian product $\mathcal{O}\times\mathbb{R}^{d}\subseteq\mathbb{R}^{d}\times\mathbb{R}^{d}\cong\mathbb{R}^{2d}$ is equipped with the Euclidean distance. By definition, ${\rm d}_{1},{\rm d}_{2}$ are metrics on $\mathbb{S}_{\mathcal{O}}=\{g:\mathcal{O}\rightrightarrows\mathbb{R}^{d}:g=\partial f~{}\text{for some locally weakly convex $f:\mathcal{O}\to\mathbb{R}$}\}$. The metric ${\rm d}_{2}$ is known as the graphical distance, aligning well with applications of Attouch’s epigraphical convergence theorem [15], see also [16, Theorem 5.1].
We document relations between ${\rm d}_{1}$ and ${\rm d}_{2}$. For every open set $\mathcal{O}$, and every pair of locally weakly convex $f_{i}:\mathcal{O}\to\mathbb{R}$ for $i=1,2$, the following bound holds from the definition:
@@
A direct verifcation gives ${\rm d}_{1}(\partial f_{1,n},\partial f_{2,n})=1$ and ${\rm d}_{2}(\partial f_{1,n},\partial f_{2,n})=1/n$ for every $n\in\mathbb{N}$.
@@
3.2 Implications to Stochastic Weakly Convex Minimizations
Theorem 1 provides key insights into the uniform convergence of subdifferentials in stochastic minimizations. Theorem 2, which builds on top of Theorem 1, shows that for stochastic minimizing objectives with certain weak convexity requirements, the uniform convergence of subdifferential sets can be effectively understood by analyzing the uniform convergence of any pair of selections from the subdifferentials of both population and empirical objectives.
To formalize our results, we begin by introducing our setup. Let $f(\cdot,\cdot):\mathbb{R}^{d}\times\Xi\to\mathbb{R}$ denote a real-valued random function. We are sampling $\xi\sim\mathbb{P}$ where $\mathbb{P}$ is supported on $\Xi$, assumed to be a Euclidean space.

### v4

@@ 2.4 Measurability Issues
To address these measurability issues, we use the concept of outer measure $\mathbb{P}^{*}$ and inner measure $\mathbb{P}_{*}$ frequently adopted in empirical process theory [34, Section 1.2-5]. Given a probability space $(\Xi,\mathscr{G},\mathbb{P})$, every subset $B\subseteq\Xi$ is assigned with an outer measure:
$\mathbb{P}^{*}(B)=\inf\{\mathbb{P}(A)\colon B\subseteq A,A\in\mathscr{G}\}.$
Notably, $\mathbb{P}^{*}$ is subadditive: $\mathbb{P}^{*}(B_{1}\cup B_{2})\leq\mathbb{P}^{*}(B_{1})+\mathbb{P}^{*}(B_{2})$, following the union bound. Correspondingly, the inner measure is $\mathbb{P}_{*}(B)=\sup\{\mathbb{P}(A)\colon B\supseteq A,A\in\mathscr{G}\}$. Clearly, $\mathbb{P}_{*}(B)=1-\mathbb{P}^{*}(\Xi\backslash B)$. Finally, when $B\in\mathscr{G}$ is measurable, then the inner and outer measure agree: $\mathbb{P}^{*}(B)=\mathbb{P}_{*}(B)=\mathbb{P}(B)$.
@@ 3.1 Uniform Subdifferential Bounds via Selections
By Rademacher’s theorem (e.g.,[17, Theorem 9.6.0]), both functions $f_{i}$ are almost everywhere differentiable in $\mathcal{O}$. Clarke’s subdifferential characterization (e.g., [17, Theorem 9.6.1]) further states that the subdifferential set $\partial f_{i}(x)$ can be expressed as:
$\begin{split}\partial f_{i}(x)&={\rm conv}\{\mathcal{G}_{i}(x)\}\end{split}$ (3.2)
where
$\mathcal{G}_{i}(x)=\left\{\lim_{k\to\infty}\nabla f_{i}(x_{k})\colon x_{k}\in D_{i},x_{k}\to x,\lim_{k\to\infty}\nabla f_{i}(x_{k})~{}\text{exists}\right\}.$
Here, $D_{i}$ denotes the set of differentiable points of $f_{i}$, which is dense in the open set $\mathcal{O}$ (meaning that $\mathcal{O}\backslash D_{i}$ has Lebesgue measure 00), and ${\rm conv}$ represents the convex hull of a set. By definition, the sets $\mathcal{G}_{i}(x)$ are closed at every $x$, and the set-valued mappings $x\mapsto\mathcal{G}_{i}(x)$ are locally bounded as the functions $f_{i}$ are locally Lipschitz.
To upper bound the Hausdorff distance between the subdifferential sets $\partial f_{1}(x)$ and $\partial f_{2}(x)$, we use a basic property of the Hausdorff distance:
$\mathbb{H}({\rm conv}\{A_{1}\},{\rm conv}\{A_{2}\})\leq\mathbb{H}(A_{1},A_{2})~{}~{}~{}\text{for any sets $A_{1}$ and $A_{2}$}.$
Applying this property to $\partial f_{1}(x)$ and $\partial f_{2}(x)$, we obtain for every $x\in\mathcal{O}$:
$\mathbb{H}(\partial f_{1}(x),\partial f_{2}(x))=\mathbb{H}({\rm conv}\{\mathcal{G}_{1}(x)\},{\rm conv}\{\mathcal{G}_{2}(x)\})\leq\mathbb{H}(\mathcal{G}_{1}(x),\mathcal{G}_{2}(x)).$ (3.3)
Therefore, the problem reduces to bounding the Hausdorff distance between $\mathcal{G}_{1}(x)$ and $\mathcal{G}_{2}(x)$. By definition, the Hausdorff distance between these sets is given by:
$\mathbb{H}(\mathcal{G}_{1}(x),\mathcal{G}_{2}(x))=\max\left\{\sup_{v_{1}\in\mathcal{G}_{1}(x)}\inf_{v_{2}\in\mathcal{G}_{2}(x)}\left\|{v_{1}-v_{2}}\right\|,\sup_{v_{2}\in\mathcal{G}_{2}(x)}\inf_{v_{1}\in\mathcal{G}_{1}(x)}\left\|{v_{1}-v_{2}}\right\|\right\}.$
Our specific goal is to bound the first term, showing that for every $x\in\mathcal{O}$:
$\sup_{v_{1}\in\mathcal{G}_{1}(x)}\inf_{v_{2}\in\mathcal{G}_{2}(x)}\left\|{v_{1}-v_{2}}\right\|\leq\sup_{x\in\mathcal{O}}\left\|{g_{1}(x)-g_{2}(x)}\right\|,$ (3.4)
since a completely symmetric argument would show the same statement holds to the second term.
Fix $x\in\mathcal{O}$. To establish (3.4), let $v_{1}^{*}\in\mathcal{G}_{1}(x)$ be the subgradient achieving the supremum, which exists due to the compactness of the set $\mathcal{G}_{1}(x)$. By definition, there exists a sequence $\{x_{k}\}$ such that $x_{k}\to x$, $x_{k}\in D_{1}$ (the set of differentiable points of $f_{1}$) and $\nabla f_{1}(x_{k})\to v_{1}^{*}$.
We now argue that we can, without loss of generality, assume that for every $k$, $x_{k}\in D_{2}$ (the set of differentiable points of $f_{2}$) as well. If some $x_{k}\not\in D_{2}$, applying Clarke’s subdifferential characterization to $f_{1}$ at $x_{k}$ (3.2), we can replace such $x_{k}$ with some $x_{k}^{\prime}$ belonging to the subset $D_{1}\cap D_{2}$ dense in $\mathcal{O}$ (note $\mathcal{O}\backslash(D_{1}\cap D_{2})$ has Lebesgue measure 00) such that $\left\|{x_{k}^{\prime}-x_{k}}\right\|\to 0$ and $\left\|{\nabla f_{1}(x_{k}^{\prime})-\nabla f_{1}(x_{k})}\right\|\to 0$ as $k\to\infty$. As a result, $x_{k}^{\prime}$ preserves all the properties of $x_{k}$, including $x_{k}^{\prime}\to x$, $x_{k}^{\prime}\in D_{1}$ and $\nabla f_{1}(x_{k}^{\prime})\to v_{1}^{*}$, while guaranteeing that $x_{k}^{\prime}\in D_{2}$ for all $k$.
Let $v_{2}^{*}$ be any accumulation point of $\nabla f_{2}(x_{k})$ (which exists since $\mathcal{G}_{2}$ are locally bounded near $x$). By taking a subsequence if necessary, we can assume that $\nabla f_{2}(x_{k})\to v_{2}^{*}$. Then,
$\sup_{v_{1}\in\mathcal{G}_{1}(x)}\inf_{v_{2}\in\mathcal{G}_{2}(x)}\left\|{v_{1}-v_{2}}\right\|\leq\left\|{v_{1}^{*}-v_{2}^{*}}\right\|=\lim_{k\to\infty}\left\|{\nabla f_{1}(x_{k})-\nabla f_{2}(x_{k})}\right\|.$ (3.5)
At differentiable points $x_{k}\in D_{1}\cap D_{2}$, the subdifferential is a singleton: $\partial f_{i}(x_{k})=\{\nabla f_{i}(x_{k})\}$, so the subgradient selection must obey $g_{i}(x_{k})=\nabla f_{i}(x_{k})$. Therefore,
$\sup_{v_{1}\in\mathcal{G}_{1}(x)}\inf_{v_{2}\in\mathcal{G}_{2}(x)}\left\|{v_{1}-v_{2}}\right\|\leq\lim_{k\to\infty}\left\|{g_{1}(x_{k})-g_{2}(x_{k})}\right\|\leq\sup_{x\in\mathcal{O}}\left\|{g_{1}(x)-g_{2}(x)}\right\|.$
This establishes (3.4). Since the argument is symmetric, we conclude:
$\mathbb{H}(\partial f_{1}(x),\partial f_{2}(x))\leq\mathbb{H}(\mathcal{G}_{1}(x),\mathcal{G}_{2}(x))\leq\sup_{x\in\mathcal{O}}\left\|{g_{1}(x)-g_{2}(x)}\right\|.$ (3.6)
Finally, since the above inequality holds for all $x\in\mathcal{O}$, this completes the proof of Theorem 1.
@@
Let $\mathcal{O}\subseteq\mathbb{R}^{d}$ be open. For a pair of locally weakly convex functions $f_{i}\colon\mathcal{O}\to\mathbb{R}$ for $i=1,2$, we define:
@@
In defining ${\rm d}_{2}$, for a locally weakly convex $f\colon\mathcal{O}\to\mathbb{R}$, we recall the graph of its subdifferential:
@@
where the Cartesian product $\mathcal{O}\times\mathbb{R}^{d}\subseteq\mathbb{R}^{d}\times\mathbb{R}^{d}\cong\mathbb{R}^{2d}$ is equipped with Euclidean distance. By definition, ${\rm d}_{1},{\rm d}_{2}$ are metrics on $\mathbb{S}_{\mathcal{O}}=\{g\colon\mathcal{O}\rightrightarrows\mathbb{R}^{d}\colon g=\partial f~{}\text{for some locally weakly convex $f\colon\mathcal{O}\to\mathbb{R}$}\}$. The metric ${\rm d}_{2}$ is known as the graphical distance, aligning well with applications of Attouch’s epigraphical convergence theorem [15], see also [16, Theorem 5.1].
We document relations between ${\rm d}_{1}$ and ${\rm d}_{2}$. For every open set $\mathcal{O}$, and every pair of locally weakly convex $f_{i}\colon\mathcal{O}\to\mathbb{R}$ for $i=1,2$, the following bound holds from the definition:
@@
A direct verification gives ${\rm d}_{1}(\partial f_{1,n},\partial f_{2,n})=1$ and ${\rm d}_{2}(\partial f_{1,n},\partial f_{2,n})=1/n$ for every $n\in\mathbb{N}$.
@@
While not essential to the rest of the paper, our proof of Theorem 1 naturally extends to the following generalization, Theorem 1, which could be of independent interest to specialists in variational analysis. To state it precisely, we recall the definition of the Clarke subdifferential of a function $f$ from the standard reference [17, equation 8(32)].
Theorem 1.
Let $f_{1}$ and $f_{2}$ be locally Lipschitz functions from $\mathcal{O}$ to $\mathbb{R}$, where $\mathcal{O}$ is an open set in $\mathbb{R}^{d}$. Let ¯$\bar{\partial}f_{1}$ and ¯$\bar{\partial}f_{2}$ denote the Clarke subdifferential of $f_{1}$ and $f_{2}$, respectively. For every pair of selections $g_{1}$ and $g_{2}$ of the subdifferentials ¯$\bar{\partial}f_{1}$ and ¯$\bar{\partial}f_{2}$ on $\mathcal{O}$, respectively, i.e., obeying $g_{1}(x)\in\bar{\partial}f_{1}(x)$ and $g_{2}(x)\in\bar{\partial}f_{2}(x)$ for all $x\in\mathcal{O}$, the following inequality holds:
$\sup_{x\in\mathcal{O}}\mathbb{H}(\bar{\partial}f_{1}(x),\bar{\partial}f_{2}(x))\leq\sup_{x\in\mathcal{O}}\left\|{g_{1}(x)-g_{2}(x)}\right\|.$ (3.10)
Proof.
By [17, Theorem 8.49 + Theorem 9.13], ¯$\bar{\partial}f_{i}(x)$ is a convex and compact set at every $x\in\mathcal{O}$. In fact, by Rademacher’s theorem [17, Theorem 9.61], ¯$\bar{\partial}f_{i}(x)$ obeys the characterization:
¯$\bar{\partial}f_{i}(x)={\rm conv}\left\{\lim_{k\to\infty}\nabla f_{i}(x_{k})\colon x_{k}\in D_{i},x_{k}\to x,\lim_{k\to\infty}\nabla f_{i}(x_{k})~{}\text{exists}\right\}.$
Given these key ingredients, the proof of Theorem 1 follows verbatim from that of Theorem 1, with the sole exception that $\partial f_{i}$ is replaced by ¯$\bar{\partial}f_{i}$. ∎
3.2 Implications to Stochastic Weakly Convex Minimization Problems
Theorem 1 provides key insights into the uniform convergence of subdifferentials in stochastic minimization problems. Theorem 2, which builds on top of Theorem 1, shows that for stochastic minimizing objectives with certain weak convexity requirements, the uniform convergence of subdifferential sets can be effectively understood by analyzing the uniform convergence of any pair of selections from the subdifferentials of both population and empirical objectives.
To formalize our results, we begin by introducing our setup. Let $f(\cdot,\cdot)\colon\mathbb{R}^{d}\times\Xi\to\mathbb{R}$ denote a real-valued random function. We are sampling $\xi\sim\mathbb{P}$ where $\mathbb{P}$ is supported on $\Xi$, assumed to be a Euclidean space.